OpenStudy (anonymous):

f(x) = -3(1.02)x x g(x) -1 -5 0 -3 1 -1 2 1 Which statement best compares the y-intercepts of f(x) and g(x)? a) The y-intercept of f(x) is equal to the y-intercept of g(x). b) The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). c)The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). d) The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (dtan5457):

Hmm if you plug x=0 to both functions, what is the y output?

OpenStudy (anonymous):

hmmmm wym

Nnesha (nnesha):

for 2nd one you have to do "NOTHING* because y-intercept is on the y-axis when x = 0 like dtan457 said so for 2nd when when x = 0 y = ???

OpenStudy (danjs):

i think the x in the first is a power, people always miss typing the ^x for some reason

OpenStudy (dtan5457):

@DanJS that explains it, i couldn't even find the answer with the regular equation

OpenStudy (dtan5457):

in that case it will look like

OpenStudy (danjs):

a^0 = 1 so f(0) = -3*1.02^0 = -3*1 = -3

OpenStudy (danjs):

The y-intercept is the value of the function when x=0. f(0) = -3*(1.02)^0 = -3 From the chart, when x=0, g(x) = -3 f(0) = g(0) = -3

OpenStudy (anonymous):

wow i am so lost

Nnesha (nnesha):

i can understand

Nnesha (nnesha):

so y-intercept always on y - axis right when x = 0 ALWAYS

OpenStudy (dtan5457):

g(x) states then when x=0, y=-3, which makes the intercept -3. -3(1.02)^0<<<<what does that equal? (that is x=0)

OpenStudy (dtan5457):

note: anything ^0=1

OpenStudy (anonymous):

it equals -3

OpenStudy (dtan5457):

exactly. if x=0, and both y=-3, what is the answer to your question?

OpenStudy (anonymous):

what is it?

Nnesha (nnesha):

what is y-intercept of f(x) and what is y-intercept of g(x) ???

OpenStudy (anonymous):

ughhhh idkkk

Nnesha (nnesha):

hmmmm did you read above comments ^^^

Nnesha (nnesha):

you just said that its equal to ??

OpenStudy (anonymous):

it equals to -3

Nnesha (nnesha):

yes f(x)<--- f of x same as y what about g(x)=y= ??

Nnesha (nnesha):

g(x) y-intercept so what is y-intercept of the table ??

OpenStudy (anonymous):

of what??? what do you mean omg

Nnesha (nnesha):

y-intercept of g(x)

OpenStudy (anonymous):

yeah but what numbers

Nnesha (nnesha):

y -intercept WHEN x equal 0 when x is 0 y is ??